Grade 11Mechanics

Two balls of equal masses are thrown upward, along the same vertical line at an interval of 2 seconds, with the same initial velocity of 14m/s. Then this collide at a height of (Take g = 10m/s2 )

Ayush kumar
8 Years agoGrade 11

1 Answer

Aryan Khera
7 Years ago
Let the displacement by ball 1 be sand displacemnt by ball 2 be s2
     Then,
             S1=14*t -1/2*10*t2
                     =14t - 5t2
          Similarly,
            S2=14(t-2) -1/2*10*(t-2)2 
As the collide,
      S1=S2
       14t -5t2=14(t-2) - 5(t-2)2
on solving we get 
       t=4.8 sec
At t=4.8 sec,
        S1=14*4.8 -5(4.8)2
           S1=48 cm as height can never be negative.